physics problem from rose

I'm having trouble with my last hw problem because I feel there is not enough information given. Please help:
A quaterback throws a football to a stationary receiver who is 31.5 m down the field. If the football is thrown at an initial angle of 40 degrees to the ground, atwhat initial speed must the quarterback throw the ball for it to reach the receiver? What is the ball's highest point during its flight?
Answer #1
the question is right

let velocity is V

therofore

Vx = Vcos(40)
Vy = Vsin(40)

use Vy to find time of flight

s = ut + .5 at^2

here s=0 , u=Vy , a= -g
so u get t
in terms of V
now 31.5 = Vx * t
solve for V (answer)

Please RATE
answered by: leanel